OpenStudy (anonymous):

(2 square root of 2 +1)(square root of 2 minus 1)

OpenStudy (anonymous):

\[(2 \sqrt{2}+1)(\sqrt{2}-1)\]

geerky42 (geerky42):

Have you ever hear of FOIL method?

OpenStudy (anonymous):

yeah, did that. got it wrong

OpenStudy (anonymous):

i got 2\[\sqrt{2}\]

geerky42 (geerky42):

You used FOIL incorrectly. Try again.